OpenStudy (anonymous):

Given that P = (2, 9) and Q = (4, 14), find the component form and magnitude of vector PQ. A) <2,5>, sqrt of 29 B) <-2,-5>, 29 C) <2,5>, 29 D) <-2,-5>, sqrt of 29 Help please!

OpenStudy (anonymous):

P=2i+9j Q=4i+14j PQ=(4-2)i+(14-9)j=2i+5j \[\left| PQ \right|=\sqrt{2^2+5^2}=?\]